# Memory-Dominated Quantum Criticality as a Universal Route to High-Temperature Superconductivity

### 16 Mar 2026 v4

@@ Appendix D Appendix D: Derivation of the Cooper-channel polarization kernel from a relaxational mode continuum
where $\xi_{\lambda}(t)$ is Gaussian white noise with
$\langle\xi_{\lambda}(t)\rangle=0,\qquad\langle\xi_{\lambda}(t)\xi_{\lambda^{\prime}}^{*}(t^{\prime})\rangle=2D_{\lambda}\,\delta(\lambda-\lambda^{\prime})\delta(t-t^{\prime}).$ (S102)
Here $D_{\lambda}$ denotes the noise strength associated with the mode $\lambda$.
In the stationary regime, obtained by sending the initial time to -$\infty$, the exact solution of Eq. (S101) is
$\Delta_{\lambda}(t)=\int_{-\infty}^{t}ds\,e^{-\lambda(t-s)}\xi_{\lambda}(s).$ (S103)
Differentiating Eq. (S103) directly yields
$\partial_{t}\Delta_{\lambda}(t)$ $=\xi_{\lambda}(t)-\lambda\int_{-\infty}^{t}ds\,e^{-\lambda(t-s)}\xi_{\lambda}(s)$
$=\xi_{\lambda}(t)-\lambda\Delta_{\lambda}(t),$ (S104)
which confirms Eq. (S101).
(II) Time-domain two-point correlation function: We now compute the stationary correlator
$C_{\lambda}(\tau)\equiv\langle\Delta_{\lambda}(\tau)\Delta_{\lambda}^{*}(0)\rangle.$ (S105)
For $\tau>0$, inserting Eq. (S103) gives
$C_{\lambda}(\tau)$ $=\int_{-\infty}^{\tau}ds\int_{-\infty}^{0}ds^{\prime}\,e^{-\lambda(\tau-s)}e^{\lambda s^{\prime}}\,\langle\xi_{\lambda}(s)\xi_{\lambda}^{*}(s^{\prime})\rangle$
$=2D_{\lambda}\int_{-\infty}^{\tau}ds\int_{-\infty}^{0}ds^{\prime}\,e^{-\lambda(\tau-s)}e^{\lambda s^{\prime}}\delta(s-s^{\prime}).$ (S106)
Because the delta function enforces $s=s^{\prime}$, while the second integral restricts $s^{\prime}\leq 0$, the effective integration domain reduces to $s\leq 0$. Therefore
$C_{\lambda}(\tau)$ $=2D_{\lambda}\int_{-\infty}^{0}ds\,e^{-\lambda(\tau-s)}e^{\lambda s}$
$=2D_{\lambda}e^{-\lambda\tau}\int_{-\infty}^{0}ds\,e^{2\lambda s}$
$=2D_{\lambda}e^{-\lambda\tau}\left[\frac{e^{2\lambda s}}{2\lambda}\right]_{-\infty}^{0}$
$=\frac{D_{\lambda}}{\lambda}e^{-\lambda\tau}.$ (S107)
By stationarity and Hermitian symmetry, the result extends to all times as
$C_{\lambda}(t)=\frac{D_{\lambda}}{\lambda}\,e^{-\lambda|t|}.$ (S108)
Equation (S108) is the exact two-point correlation function of a single Ornstein–Uhlenbeck pair mode. The factor $1/\lambda$ in front of the exponential is crucial. It is precisely this factor that removes the extra numerator $\lambda$ appearing in the Fourier transform of the bare retarded kernel.
(III) Frequency-domain spectral density: The mode-resolved fluctuation spectrum is defined as
$S_{\lambda}(\omega)\equiv\int_{-\infty}^{\infty}dt\,e^{i\omega t}C_{\lambda}(t).$ (S109)
Substituting Eq. (S108) gives
$S_{\lambda}(\omega)=\frac{D_{\lambda}}{\lambda}\int_{-\infty}^{\infty}dt\,e^{i\omega t}e^{-\lambda|t|}.$ (S110)
Splitting the integral into positive and negative times,
$S_{\lambda}(\omega)$ $=\frac{D_{\lambda}}{\lambda}\left[\int_{0}^{\infty}dt\,e^{-(\lambda-i\omega)t}+\int_{-\infty}^{0}dt\,e^{(\lambda+i\omega)t}\right].$ (S111)
Changing variables u=tu=-t in the second term yields
$\int_{-\infty}^{0}dt\,e^{(\lambda+i\omega)t}=\int_{0}^{\infty}du\,e^{-(\lambda+i\omega)u},$ (S112)
so that
$S_{\lambda}(\omega)=\frac{D_{\lambda}}{\lambda}\left[\int_{0}^{\infty}dt\,e^{-(\lambda-i\omega)t}+\int_{0}^{\infty}dt\,e^{-(\lambda+i\omega)t}\right].$ (S113)
@@
$\int_{0}^{\infty}dt\,e^{-at}=\frac{1}{a},\qquad\Re a>0,$ (S114)
we obtain
$S_{\lambda}(\omega)$ $=\frac{D_{\lambda}}{\lambda}\left(\frac{1}{\lambda-i\omega}+\frac{1}{\lambda+i\omega}\right)$
$=\frac{D_{\lambda}}{\lambda}\frac{2\lambda}{\lambda^{2}+\omega^{2}}.$ (S115)
@@
$S_{\lambda}(\omega)=\frac{2D_{\lambda}}{\lambda^{2}+\omega^{2}}.$ (S116)
This is the key result. The Fourier transform of the exponential memory kernel itself produces a numerator $\lambda$, but the stationary correlator contains a compensating factor $1/\lambda$. As a result, the exact fluctuation spectrum carries no additional factor of $\lambda$ in the numerator.
(IV) Equivalent derivation in frequency space: The same result follows directly in frequency space. Fourier transforming Eq. (S101) gives
$(-i\omega+\lambda)\Delta_{\lambda}(\omega)=\xi_{\lambda}(\omega),$ (S117)
hence
$\Delta_{\lambda}(\omega)=\frac{1}{\lambda-i\omega}\,\xi_{\lambda}(\omega).$ (S118)
The noise correlator in frequency space is
$\langle\xi_{\lambda}(\omega)\xi_{\lambda^{\prime}}^{*}(\omega^{\prime})\rangle=2\pi\,2D_{\lambda}\,\delta(\lambda-\lambda^{\prime})\delta(\omega-\omega^{\prime}).$ (S119)
Therefore
$\langle\Delta_{\lambda}(\omega)\Delta_{\lambda^{\prime}}^{*}(\omega^{\prime})\rangle$ $=\frac{\langle\xi_{\lambda}(\omega)\xi_{\lambda^{\prime}}^{*}(\omega^{\prime})\rangle}{(\lambda-i\omega)(\lambda^{\prime}+i\omega^{\prime})}$
$=2\pi\,\delta(\lambda-\lambda^{\prime})\delta(\omega-\omega^{\prime})\frac{2D_{\lambda}}{\lambda^{2}+\omega^{2}}.$ (S120)
Defining the spectral density through
$\langle\Delta_{\lambda}(\omega)\Delta_{\lambda}^{*}(\omega^{\prime})\rangle=2\pi\delta(\omega-\omega^{\prime})S_{\lambda}(\omega),$ (S121)
one again recovers Eq. (S116).
This derivation makes clear that the mode-resolved pair correlator is proportional to $D_{\lambda}^{R}(\omega)D_{\lambda}^{A}(\omega)$, not to the real part of a single retarded propagator. Indeed,
$D_{\lambda}^{R}(\omega)=\frac{1}{\lambda-i\omega},\qquad D_{\lambda}^{A}(\omega)=\frac{1}{\lambda+i\omega},$ (S122)
so that
$D_{\lambda}^{R}(\omega)D_{\lambda}^{A}(\omega)=\frac{1}{\lambda^{2}+\omega^{2}}.$ (S123)
(V) From a single mode to the Cooper-channel polarization: After microscopic electronic degrees of freedom are integrated out, the collective Cooper sector is represented as a continuum of relaxational pair modes. Let $\rho_{\rm pair}(\lambda)$ denote the effective density of pair relaxation rates, including both mode degeneracy and noise weight. The full polarization kernel is then obtained by summing the mode-resolved fluctuation spectrum over the continuum,
$\Pi(\omega)=\int_{0}^{\Lambda}d\lambda\,\rho_{\rm pair}(\lambda)\,S_{\lambda}(\omega).$ (S124)
Absorbing the factor $2D_{\lambda}$ into the definition of $\rho_{\rm pair}(\lambda)$ gives
$\Pi(\omega)=\int_{0}^{\Lambda}d\lambda\,\rho_{\rm pair}(\lambda)\,\frac{1}{\lambda^{2}+\omega^{2}}.$ (S125)
This is the Cooper-channel polarization kernel quoted in the main text.
(VI) Infrared scaling of the polarization kernel: For a generic infrared TDOS
$\rho_{\rm pair}(\lambda)\sim\lambda^{\eta},\qquad\lambda\to 0,$ (S126)
Eq. (S125) gives
$\Pi(\omega)\sim\int_{0}^{\Lambda}d\lambda\,\frac{\lambda^{\eta}}{\lambda^{2}+\omega^{2}}.$ (S127)
Setting $\lambda=|\omega|x$,
$\Pi(\omega)$ $\sim|\omega|^{\eta-1}\int_{0}^{\Lambda/|\omega|}dx\,\frac{x^{\eta}}{1+x^{2}}.$ (S128)
Hence the universal infrared scaling is
$\Pi(\omega)\sim|\omega|^{\eta-1}.$ (S129)
In particular, for a flat TDOS,
$\rho_{\rm pair}(\lambda\to 0)=\rho_{0},$ (S130)
@@
$\Pi(\omega)$ $\simeq\rho_{0}\int_{0}^{\Lambda}\frac{d\lambda}{\lambda^{2}+\omega^{2}}$
$=\frac{\rho_{0}}{|\omega|}\tan^{-1}\!\left(\frac{\Lambda}{|\omega|}\right)$
$\xrightarrow[|\omega|\ll\Lambda]{}\frac{\pi\rho_{0}}{2}\frac{1}{|\omega|}.$ (S131)
Equation (S131) establishes the memory-dominated Cooper enhancement. A flat relaxation-rate spectrum therefore converts the marginal logarithmic pairing growth of conventional Ohmic dynamics into an algebraically enhanced infrared instability of the Cooper channel.
Appendix E Appendix E: Spectral function generated by memory–dominated self–energy
where
Writing $\chi_{R}^{-1}=X(\omega,\mathbf{q})+iY(\omega)$ with
$Y(\omega)$ $=\Gamma\,\mathrm{sgn}(\omega),$ (S137)
one has $\chi_{R}=(X-iY)/(X^{2}+Y^{2})$, hence
which yields
(V) Universal scaling collapse: Define the scaling variable
Using $\Gamma=(\pi/2)\rho_{0}$, Eq. (S138) reduces to the universal form
(VI) Finite–temperature infrared cutoff: At finite temperature the logarithmic growth is regulated by the thermal scale. A minimal phenomenological implementation consistent with Appendix D is
$|\omega|\;\rightarrow\;\max\{|\omega|,cT\},\qquad c=O(1),$ (S145)
yielding
Appendix F Appendix F: Possible connections between flat TDOS and scale–free dynamical phenomena

### 30 Mar 2026 v5

@@ Appendix C Appendix C: Exact emergence of a memory kernel from integrating out an OU reservoir
where $\xi_{\lambda}(t)$ is a stochastic source. The deterministic linear-response part is
$(\partial_{t}+\lambda)\Delta_{\lambda}(t)=\xi_{\lambda}(t).$ (S107)
The retarded Green’s function $L_{\lambda}^{R}(t-t^{\prime})$ is defined as the inverse of the differential operator $\partial_{t}+\lambda$, namely
[…]
For $t\neq t^{\prime}$, the delta function vanishes and Eq. (S108) reduces to
$(\partial_{t}+\lambda)L_{\lambda}^{R}(t-t^{\prime})=0.$ (S109)
Hence
$L_{\lambda}^{R}(t-t^{\prime})=A\,e^{-\lambda(t-t^{\prime})}\quad\text{for }t>t^{\prime}.$ (S110)
Integrating Eq. (S108) across $t=t^{\prime}$ gives
$\int_{t^{\prime}-\epsilon}^{t^{\prime}+\epsilon}dt\,(\partial_{t}+\lambda)L_{\lambda}^{R}(t-t^{\prime})=\int_{t^{\prime}-\epsilon}^{t^{\prime}+\epsilon}dt\,\delta(t-t^{\prime})=1.$ (S111)
Since the integral of the regular term $\lambda L_{\lambda}^{R}$ vanishes as $\epsilon\to 0$, one obtains
$L_{\lambda}^{R}(t^{\prime}+0)-L_{\lambda}^{R}(t^{\prime}-0)=1.$ (S112)
With the retarded boundary condition $L_{\lambda}^{R}(t^{\prime}-0)=0$, this yields $A=1$, so that
$L_{\lambda}^{R}(t-t^{\prime})=\theta(t-t^{\prime})\,e^{-\lambda(t-t^{\prime})}.$ (S113)
Fourier transforming,
$L_{\lambda}^{R}(\omega)=\int_{-\infty}^{\infty}dt\,e^{i\omega t}\theta(t)e^{-\lambda t}=\int_{0}^{\infty}dt\,e^{-(\lambda-i\omega)t}=\frac{1}{\lambda-i\omega}.$ (S114)
Similarly,
$L_{\lambda}^{A}(\omega)=\frac{1}{\lambda+i\omega}.$ (S115)
Thus a single relaxational pair mode contributes the retarded response kernel
$L_{\lambda}^{R}(\omega)=\frac{1}{\lambda-i\omega}.$ (S116)
(III) Continuum of relaxational pair modes and TDOS: We now assume that, after integrating out microscopic fermionic degrees of freedom, the low-energy Cooper sector is represented by a continuum of relaxational pair modes labeled by $\lambda$. Let $\rho_{\rm pair}(\lambda)$ denote the TDOS of pair relaxation rates.
The irreducible retarded Cooper kernel is then obtained by summing the mode-resolved retarded response over the continuum:
$\Pi^{R}(\omega;T)=\int_{T}^{\Lambda}d\lambda\,\rho_{\rm pair}(\lambda)\,\frac{1}{\lambda-i\omega}.$ (S117)
Here $\Lambda$ is the ultraviolet cutoff of the collective theory, and the lower limit is taken to be $T$, which serves as the infrared thermal cutoff in the static instability problem.
Equation (S117) is the retarded analogue of the usual Cooper bubble after projection onto the collective pair sector. It is this $\Pi^{R}$, rather than the Keldysh fluctuation correlator, that enters the linear-response instability criterion.
It is useful to separate real and imaginary parts:
$\Pi^{R}(\omega;T)$ $=\int_{T}^{\Lambda}d\lambda\,\rho_{\rm pair}(\lambda)\,\frac{\lambda+i\omega}{\lambda^{2}+\omega^{2}}$
$=\int_{T}^{\Lambda}d\lambda\,\rho_{\rm pair}(\lambda)\,\frac{\lambda}{\lambda^{2}+\omega^{2}}$
$\quad+i\omega\int_{T}^{\Lambda}d\lambda\,\rho_{\rm pair}(\lambda)\,\frac{1}{\lambda^{2}+\omega^{2}}.$ (S118)
Therefore
$\Re\Pi^{R}(\omega;T)=\int_{T}^{\Lambda}d\lambda\,\rho_{\rm pair}(\lambda)\,\frac{\lambda}{\lambda^{2}+\omega^{2}},$ (S119)
and
$\Im\Pi^{R}(\omega;T)=\omega\int_{T}^{\Lambda}d\lambda\,\rho_{\rm pair}(\lambda)\,\frac{1}{\lambda^{2}+\omega^{2}}.$ (S120)
In the static limit $\omega\to 0$, one obtains
$\Pi^{R}(0;T)=\int_{T}^{\Lambda}d\lambda\,\frac{\rho_{\rm pair}(\lambda)}{\lambda}.$ (S121)
This is the quantity that controls the transition temperature.
(IV) Retarded collective ladder and Thouless criterion: Let $g_{\rm pair}>0$ denote the short-distance irreducible pairing vertex in the projected Cooper sector. The retarded ladder equation for the full pair susceptibility is
$\chi_{\Delta}^{R}(\omega;T)=\Pi^{R}(\omega;T)+\Pi^{R}(\omega;T)\,g_{\rm pair}\,\chi_{\Delta}^{R}(\omega;T).$ (S122)
Solving Eq. (S122) for $\chi_{\Delta}^{R}$, we obtain
$\chi_{\Delta}^{R}(\omega;T)=\frac{\Pi^{R}(\omega;T)}{1-g_{\rm pair}\Pi^{R}(\omega;T)}.$ (S123)
Equivalently, the series form is
$\chi_{\Delta}^{R}(\omega;T)=$ $\Pi^{R}(\omega;T)+\Pi^{R}(\omega;T)g_{\rm pair}\Pi^{R}(\omega;T)$
$+\Pi^{R}(\omega;T)g_{\rm pair}\Pi^{R}(\omega;T)g_{\rm pair}\Pi^{R}(\omega;T)+\cdots.$ (S124)
The superconducting instability occurs when the retarded susceptibility develops a pole at zero external frequency and momentum. Therefore the Thouless criterion is
$1-g_{\rm pair}\Pi^{R}(0;T_{c})=0.$ (S125)
Using Eq. (S121), the transition scale is determined by
$1-g_{\rm pair}\int_{T_{c}}^{\Lambda}d\lambda\,\frac{\rho_{\rm pair}(\lambda)}{\lambda}=0.$ (S126)
Equation (S126) is the central formula for the transition temperature in the retarded linear-response formulation.
(V) General infrared TDOS exponent $\alpha$: We now assume the infrared TDOS
$\rho_{\rm pair}(\lambda)=\rho_{\alpha}\,\lambda^{\alpha},\qquad T\leq\lambda\leq\Lambda,$ (S127)
with $\rho_{\alpha}>0$. Then Eq. (S121) becomes
$\Pi^{R}(0;T)=\rho_{\alpha}\int_{T}^{\Lambda}d\lambda\,\lambda^{\alpha-1}.$ (S128)
For $\alpha\neq 0$,
$\int d\lambda\,\lambda^{\alpha-1}=\frac{\lambda^{\alpha}}{\alpha},$ (S129)
hence
$\Pi^{R}(0;T)=\frac{\rho_{\alpha}}{\alpha}\left(\Lambda^{\alpha}-T^{\alpha}\right),\qquad\alpha\neq 0.$ (S130)
For $\alpha=0$,
$\Pi^{R}(0;T)=\rho_{0}\int_{T}^{\Lambda}\frac{d\lambda}{\lambda}=\rho_{0}\ln\!\left(\frac{\Lambda}{T}\right).$ (S131)
Substituting into the Thouless criterion $1-g_{\rm pair}\Pi^{R}(0;T_{c})=0$, one finds
$1-g_{\rm pair}\frac{\rho_{\alpha}}{\alpha}\left(\Lambda^{\alpha}-T_{c}^{\alpha}\right)=0,\qquad\alpha\neq 0,$ (S132)
and
$1-g_{\rm pair}\rho_{0}\ln\!\left(\frac{\Lambda}{T_{c}}\right)=0,\qquad\alpha=0.$ (S133)
It is useful to separate three regimes.
(a) $\alpha>0$:
$\Pi^{R}(0;T\to 0)\to\frac{\rho_{\alpha}}{\alpha}\Lambda^{\alpha},$ (S134)
which is finite. Therefore there is no automatic infrared divergence in the Cooper kernel. A superconducting instability exists only if
$g_{\rm pair}\frac{\rho_{\alpha}}{\alpha}\Lambda^{\alpha}>1.$ (S135)
(b) $\alpha=0$: the kernel diverges logarithmically and yields a BCS-like transition scale.
(c) $\alpha<0$: the kernel diverges algebraically as $T\to 0$, producing an IR-singular enhancement of pairing.
(VI) Case $\alpha=0$ (flat TDOS and BCS-like exponential scale): For a flat infrared TDOS,
$\rho_{\rm pair}(\lambda)=\rho_{0},$ (S136)
Eq. (S131) gives
$\Pi^{R}(0;T)=\rho_{0}\ln\!\left(\frac{\Lambda}{T}\right).$ (S137)
The Thouless criterion results in
$T_{c}=\Lambda\exp\!\left(-\frac{1}{g_{\rm pair}\rho_{0}}\right).$ (S138)
Thus a flat TDOS leads to a BCS-like exponential transition scale. The physical interpretation, however, is different from weak-coupling BCS: the logarithm is controlled not by the quasiparticle density of states, but by the infrared density of slow collective pair relaxation modes.
(VII) Case $\alpha=-1/2$ (IR-singular algebraic enhancement): We now consider the singular infrared TDOS
$\rho_{\rm pair}(\lambda)=\rho_{-1/2}\lambda^{-1/2}.$ (S139)
Then
$\Pi^{R}(0;T)=\rho_{-1/2}\int_{T}^{\Lambda}d\lambda\,\lambda^{-3/2}.$ (S140)
@@
$\Pi^{R}(0;T)$ $=\rho_{-1/2}\left[-2\lambda^{-1/2}\right]_{T}^{\Lambda}$
$=2\rho_{-1/2}\left(T^{-1/2}-\Lambda^{-1/2}\right),$ (S141)
the Thouless criterion becomes
$1-2g_{\rm pair}\rho_{-1/2}\left(T_{c}^{-1/2}-\Lambda^{-1/2}\right)=0.$ (S142)
The exact transition scale is
$T_{c}=\left(\Lambda^{-1/2}+\frac{1}{2g_{\rm pair}\rho_{-1/2}}\right)^{-2}.$ (S143)
In the weak-coupling limit
$\frac{1}{2g_{\rm pair}\rho_{-1/2}}\gg\Lambda^{-1/2},$ (S144)
the ultraviolet term may be neglected, so that
$T_{c}\simeq\left(2g_{\rm pair}\rho_{-1/2}\right)^{2}.$ (S145)
Thus the $\alpha=-1/2$ IR-singular TDOS converts the BCS-like exponential scale into an algebraically amplified transition scale.
(VIII) General algebraic regime for $\alpha<0$: For completeness, let $\alpha=-\beta$ with $\beta>0$. Then Eq. (S130) becomes
$\Pi^{R}(0;T)=\frac{\rho_{-\beta}}{\beta}\left(T^{-\beta}-\Lambda^{-\beta}\right).$ (S146)
The Thouless criterion yields
$1-g_{\rm pair}\frac{\rho_{-\beta}}{\beta}\left(T_{c}^{-\beta}-\Lambda^{-\beta}\right)=0,$ (S147)
@@
$T_{c}=\left(\Lambda^{-\beta}+\frac{\beta}{g_{\rm pair}\rho_{-\beta}}\right)^{-1/\beta}.$ (S148)
In the weak-coupling infrared-dominated regime,
$T_{c}\simeq\left(\frac{g_{\rm pair}\rho_{-\beta}}{\beta}\right)^{1/\beta}.$ (S149)
Thus the retarded Keldysh formulation yields the following general result:
$\alpha=0\;\Rightarrow\;T_{c}\sim\Lambda e^{-1/(g_{\rm pair}\rho_{0})},\quad\alpha<0\;\Rightarrow\;T_{c}\sim g_{\rm pair}^{\,1/|\alpha|}.$ (S150)
Final remark. The derivation above concerns only the retarded linear-response criterion for the superconducting instability. The Keldysh fluctuation propagator,
$L^{K}(\omega)=L^{R}(\omega)\Sigma^{K}(\omega)L^{A}(\omega),$ (S151)
and the corresponding fluctuation correlator will be analyzed separately in the subsequent part of this Appendix. The essential logical point is that the transition scale $T_{c}$ is fixed first by the pole of the retarded ladder, while the fluctuation sector is discussed afterwards.
(IX) Keldysh fluctuation propagator of the collective pair field: Having determined the transition scale from the retarded linear-response kernel, we now briefly summarize the corresponding fluctuation sector. In the Keldysh formalism, the fluctuation propagator is obtained from the standard matrix relation
$L^{K}(\omega)=L^{R}(\omega)\,\Sigma^{K}(\omega)\,L^{A}(\omega),$ (S152)
where $\Sigma^{K}(\omega)$ is the Keldysh noise kernel. For the present purpose, the precise microscopic form of $\Sigma^{K}$ is not essential; what matters is that $L^{K}$ is built from the same retarded and advanced collective propagators that control the instability criterion.
For a single relaxational pair mode with rate $\lambda$,
$L_{\lambda}^{R}(\omega)=\frac{1}{\lambda-i\omega},\qquad L_{\lambda}^{A}(\omega)=\frac{1}{\lambda+i\omega}.$ (S153)
If the corresponding Keldysh noise strength is denoted by $\Sigma_{\lambda}^{K}(\omega)$, then
$L_{\lambda}^{K}(\omega)=L_{\lambda}^{R}(\omega)\,\Sigma_{\lambda}^{K}(\omega)\,L_{\lambda}^{A}(\omega)=\frac{\Sigma_{\lambda}^{K}(\omega)}{\lambda^{2}+\omega^{2}}.$ (S154)
In the simplest low-frequency approximation, one may treat $\Sigma_{\lambda}^{K}(\omega)$ as slowly varying and replace it by an effective constant amplitude. Then the essential frequency dependence is Lorentzian,
$L_{\lambda}^{K}(\omega)\propto\frac{1}{\lambda^{2}+\omega^{2}}.$ (S155)
(X) Continuum TDOS representation of the fluctuation kernel: Summing over the continuum of collective pair modes, the Keldysh fluctuation kernel takes the form
$\Pi^{K}(\omega)=\int_{0}^{\Lambda}d\lambda\,\rho_{\rm pair}(\lambda)\,L_{\lambda}^{K}(\omega).$ (S156)
In the low-frequency regime, using Eq. (S155), this reduces to
$\Pi^{K}(\omega)\propto\int_{0}^{\Lambda}d\lambda\,\frac{\rho_{\rm pair}(\lambda)}{\lambda^{2}+\omega^{2}}.$ (S157)
For the infrared TDOS
$\rho_{\rm pair}(\lambda)=\rho_{\alpha}\lambda^{\alpha},$ (S158)
@@
$\Pi^{K}(\omega)\propto\rho_{\alpha}\int_{0}^{\Lambda}d\lambda\,\frac{\lambda^{\alpha}}{\lambda^{2}+\omega^{2}}.$ (S159)
Rescaling $\lambda=|\omega|x$, one finds
$\Pi^{K}(\omega)$ $\propto\rho_{\alpha}\int_{0}^{\Lambda/|\omega|}d(|\omega|x)\,\frac{(|\omega|x)^{\alpha}}{|\omega|^{2}(1+x^{2})}$
$=\rho_{\alpha}\,|\omega|^{\alpha-1}\int_{0}^{\Lambda/|\omega|}dx\,\frac{x^{\alpha}}{1+x^{2}}.$ (S160)
Hence, up to an $\alpha$-dependent numerical factor,
$\Pi^{K}(\omega)\propto|\omega|^{\alpha-1}.$ (S161)
(XI) Representative infrared cases: The fluctuation kernel therefore exhibits the following infrared behavior.
(a) $\alpha=0$: flat TDOS
$\Pi^{K}(\omega)\propto\frac{1}{|\omega|}.$ (S162)
Thus a flat TDOS produces a strongly enhanced algebraic fluctuation spectrum even though the transition scale itself remains BCS-like, $T_{c}=\Lambda\exp\!\left(-\frac{1}{g_{\rm pair}\rho_{0}}\right)$.
(b) $\alpha=-1/2$: supercritical TDOS
$\Pi^{K}(\omega)\propto|\omega|^{-3/2}.$ (S163)
The fluctuation enhancement is now even stronger than in the flat case, consistent with the algebraically amplified transition scale derived above.
(c) general $\alpha<0$:
$\Pi^{K}(\omega)\propto|\omega|^{-1-|\alpha|}.$ (S164)
showing that increasingly singular infrared TDOS leads to progressively stronger fluctuation enhancement.
(XII) Physical interpretation: The distinction between the retarded and Keldysh sectors should now be clear. The superconducting transition temperature is fixed by the pole of the retarded ladder, $1-g_{\rm pair}\Pi^{R}(0;T_{c})=0$, whereas the Keldysh kernel in Eq. (S157) describes the strength of collective Cooper-pair fluctuations. Accordingly, a flat TDOS already produces strong algebraic infrared fluctuations,
$\Pi^{K}(\omega)\propto|\omega|^{-1},$ (S165)
while a singular TDOS with $\alpha<0$ enhances both the fluctuation spectrum and, through the retarded kernel, the transition scale itself.
In this sense, the fluctuation sector and the instability sector are closely related but logically distinct: the retarded component determines the transition point, whereas the Keldysh component quantifies the infrared strength of pair-field fluctuations near that transition.
Appendix E Appendix E: Possible connections between flat TDOS and scale–free dynamical phenomena